OpenStudy (anonymous):

Find the GCF of 33 and 99 a.37 b.33 c.66 d.11

OpenStudy (anonymous):

33

OpenStudy (anonymous):

33= 3*11 and 99=3*3*11 therfore in common they have 3*11 which is 33

jhonyy9 (jhonyy9):

b. is right